OpenStudy (anonymous):

Algebra 2 help? Writing function compositions and inverses!! WILL GIVE MEDAL

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 2 }x+1\] a) find the inverse of f(x) and name it g(x). Show and explain your work. b) Use function composition to show that f(x) and g(x) are inverses of each other.

OpenStudy (anonymous):

i just need help with B now.

OpenStudy (anonymous):

@alprincenofl ^^

OpenStudy (cwrw238):

its best to do this using an example let f(x) = 3x - 2 let y = 3x - 2 3x = y + 2 x = (y = 2)/3 f'(x) = ( x + 2)/3 = g(x) if g(x) is an inverse of f(x) then f(gx)) = x so lets find f(g(x)) this = 3 ( x + 2)/3 - 2 = x +2 -2 = x this verifies that g(x) is the inverse of f(x)

OpenStudy (anonymous):

how would I do it with the fraction though?? .-. that's where I'm confused

OpenStudy (cwrw238):